A researcher wishes to estimate the mean amount of money spent per month on food by households in a certain neighborhood. She desires a margin of error of $21 . Past studies suggest that a population standard deviation of $220 is reasonable. Estimate the minimum sample size needed to estimate the population mean with the stated accuracy
Answer by Boreal(15235)

if pop sd known, use z.
half-interval of error is z*SE or z*sd/sqrt(n)
so 21=1.96*220/sqrt(n), assuming a 95% CI interval.
square both sides
441=3.84*48400/n
n=384*48400/441
=421.62 or 422 ANSWER